# From Parameters to Behaviors: Unsupervised Compression of the Policy Space

### v1

@@ 4 Method: Unsupervised Compression of the Policy Space
To address the sample inefficiency inherent in high-dimensional policy parameter spaces, we propose a paradigm shift from directly optimizing in the parameter space to learning within a compact, low-dimensional policy manifold that captures the true diversity of behaviors. This is achieved through a two-phase framework: a completely unsupervised, task-agnostic pre-training phase to discover the manifold, followed by a supervised, task-specific fine-tuning phase. As illustrated in Fig. 2, this framework is realized in three steps: $\mathbf{(1)}$ generating a behaviorally diverse dataset of policies, $\mathbf{(2)}$ learning the latent policy manifold via a generative model, and $\mathbf{(3)}$ performing fine-tuning by optimizing over this learned latent space.
Policy Dataset Generation.  Essential to many manifold reconstruction algorithms is the requirement that one can efficiently cover the manifold with samples (Bernstein et al., 2000; Cheng et al., 2005; Fefferman et al., 2016). Thus, the first stage involves generating a dataset ${\mathcal{D}}_{\Theta}$ of policies that are possibly covering the manifold of behaviorally diverse policies.
A naïve option is to randmly sample $N$ policies, $\hat{\mathcal{D}}_{\Theta}=\{{\bm{\theta}}_{i}\}_{i=1}^{N}$, by drawing their parameters from a uniform distribution. Unfortunately, it is well-understood that such naïve sampling from the parameter space is unlikely to produce uniform coverage of the behavior space, as it tends to favor functionally similar, often non-exploring, policies.
To address this bias, an explicit measure of behavioral diversity is needed. Looking at Eq. 4, one notices that optimizing such a measure directly requires to estimate the divergence between two state-action distributions $d_{\pi_{\bm{\theta}}}^{sa},d_{\pi_{{\bm{\theta}}^{\prime}}}^{sa}$. Unfortunately, this would not only be computationally intensive, but it would also require sampling from the environment for a potentially very large set of policies. To avoid this, we will take into account an upper bound to this quantity in the case of finite-horizon tasks (Prop. E.1, Metelli et al., 2018), namely $D(\pi_{\bm{\theta}}||\pi_{{\bm{\theta}}^{\prime}})$. In practice, we substitute this measure with the L2 distance of two policies in the action space, evaluated on a finite subset of the state space, or formally:
$D(\pi_{\bm{\theta}}||\pi_{{\bm{\theta}}^{\prime}})=\sqrt{\sum_{i=1}^{M}(\pi_{\bm{\theta}}(\cdot|s_{i})-\pi_{{\bm{\theta}}^{\prime}}(\cdot|s_{i}))^{2}}.\vskip-2.84544pt$ (6)
Based on this proxy, we apply a novelty search algorithm (Lehman & Stanley, 2011) by computing a novelty score, $\rho(\pi_{\bm{\theta}})$, for each policy based on its average divergence from its $k$-nearest neighbors: $\rho(\pi_{\bm{\theta}})=\frac{1}{k}\sum_{i=1}^{k}D(\pi_{\bm{\theta}}||\pi_{{\bm{\theta}}_{i}}).$
@@
Latent Behavior Compression.  In the second stage, we learn the low-dimensional manifold from the filtered policy dataset ${\mathcal{D}}_{\Theta}$. Potentially, any generative model would do the work, but here we are interested in learning latent low-dimensional representations while preserving the end-to-end differentiability that makes gradient-based optimization effective. For these reasons, we employ a symmetric autoencoder architecture with an encoder $f_{\bm{\xi}}:\Theta\to{\mathcal{Z}}$ and a decoder $g_{\bm{\zeta}}:{\mathcal{Z}}\to\Theta$. While a standard autoencoder minimizes parameter reconstruction error, our goal is to compress policy behavior. We therefore introduce a novel Behavioral Reconstruction Loss, which trains the autoencoder to minimize the expected behavioral divergence between the original policy and its reconstruction:
${\mathcal{L}}_{B}({\bm{\xi}},{\bm{\zeta}})=\mathbb{E}_{{\bm{\theta}}\sim{\mathcal{D}}_{\Theta}}\left[D\left(\pi_{\bm{\theta}}||\pi_{g_{\bm{\zeta}}(f_{\bm{\xi}}({\bm{\theta}}))}\right)\right].\vskip-5.69046pt$ (7)
This objective frees the decoder from reproducing the exact parameter values, allowing it to discover any parameterization that generates the desired behavior. As a result, the latent space $\mathcal{Z}$} becomes organized purely by functional similarity, effectively capturing the policy manifold. In practice, we use an empirical estimator of the behavioral reconstruction loss based on the notion of divergence in the action space. For this purpose, we train our autoencoders to minimize the Mean Squared Error between action vectors relative to a subset of the state space sampled at each gradient step, resulting in the estimator $\hat{\mathcal{L}}_{B}({\bm{\xi}},{\bm{\zeta}})=\frac{1}{NM}\sum_{i,j=1}^{N,M}(\pi_{{\bm{\theta}}_{i}}(s_{j})-\pi_{g_{\bm{\zeta}}(f_{\bm{\xi}}({\bm{\theta}}_{i}))}(s_{j}))^{2}$, where $N$ is the number of policies, and $M$ is the number of sampled states.
@@
$\nabla_{\bm{z}}J^{R}({\bm{z}})=\nabla_{\bm{z}}g_{{\bm{\zeta}}^{\star}}({\bm{z}})^{\top}\nabla_{\bm{\theta}}J^{R}(\theta),\vskip-2.0pt$ (8)
@@
Remarks.  In this section, we proposed three specific instantiations for each phase. Yet, we stress the fact that the proposed pipeline represents the most relevant contribution per se, independently of how it is realized, i.e., how the policies are collected, which divergence measure is used, which generative model or PO algorithm over the latent space is employed.
@@ 5 Experiments
We now investigate through extensive empirical corroboration how the proposed method addresses the research questions. In order to do so, we will mainly focus on the unsupervised pre-training phase of the proposed pipeline as of Fig. 2, in which a latent representation is built out of general datasets of policies not designed to address any specific task explicitly, and we report the empirical results in Subsec. 5.1. Finally, we make sure that such a latent space can indeed be leveraged in later supervised fine-tuning phases as soon as a task is provided, and report the results in Subsec. 5.2.
Experimental Domains.  The experiments are performed to illustrate essential features of Latent Behavior Compression, and for this reason, the domains are selected for being challenging while keeping high interpretability. The first is the Mountain Car Continuous (MC, Moore, 1990) environment. To evaluate the quality and characteristics of the latent space, we define four downstream tasks: standard and left have the goal state on the right and left hill, respectively; speed and height incentivize the car to keep a high speed and vertical coordinate, respectively, without terminating the episode. The second is the Reacher environment from the MuJoCo suite (RC, Tassa et al., 2018). For this environment, we define four downstream tasks in this environment: speed, which incentivizes the fingertip to move with high linear velocity; clockwise and c-clockwise reward the agent for each step the agent is rotating clockwise and counterclockwise, respectively; and radial, which promotes the retraction and extension of the arm. A detailed description of the environments can be found in Appendix B.
Experimental Regimes.  The experiments are performed over a set of different parameters. In MC, we took into account three Policy Sizes (Small, Medium, and Large) with roughly $10^{1},10^{3}$, and $10^{5}$ parameters respectively, three Policy Dataset Sizes (10k, 50k, and 100k generated policies, with a 10% novelty-based cut-down), and three Latent Dimensions (1D, 2D, 3D). In RC, we focused on a specific configuration with Policy Sizes medium, Policy Dataset Sizes of 100k, but five possible latent dimensions (1D, 2D, 3D, 5D, 8D). [^3] Notably, the AE architecture and training hyper-parameters are left the same for every experiment, regardless of configuration or environment.
@@ 5.1 Unsupervised Pre-training
(Q1) Is it possible to learn a low-dimensional latent representation of high-dimensional policy parameter spaces in an unsupervised fashion? (Q2) What are the properties of such a latent representation, if it exists? Is its intrinsic dimension a function of the behavioral complexity rather than the size of the parameter space?
@@
Landscape of the Latent Behavior Manifold.  We visually inspected the latent spaces trained under different conditions and environments, and we report a handful in Fig. 3. Interestingly, it is apparent that the latent spaces, regardless of the choice of D (top-to-bottom) or policy size (left-to-right), are able to encode some behaviorally diverse policies with high performance. For instance, in Fig. 3(e),3(f), and 3(g), a 2D latent space is able to encode policies of all three sizes, but the landscape grows more complex with the larger policy sizes. We speculate that this is due to the increased range of behaviors expressed by larger policies and the hardness of high-compression regimes. Indeed, the same trend is present for different tasks, as in Fig. 3(i),3(k). On the other hand, by changing the encoding dimension as in Fig. 3(b),3(f),3(j), it is clear how certain behavioral areas at high performance are able to grow larger, creating more nuanced decoded policies. Unfortunately yet, the compression is only as good as the dataset used to learn the latent space: when a behavior is scarcely represented in the dataset, as is the case for the task height in Fig. 3(a),3(c), it is unlikely that the learned representation will encode it in large areas, or encode it at all. As for RC, the compression architecture struggles to compress the policies at higher compression regimes (Fig. 3(d),3(h)), as the environment is more challenging and presents a wider range of behaviors. On the other hand, large areas of good quality compression are present for larger dimensions of the encoding (Fig. 3(l)), confirming the expected theoretical behavior (Hurewicz & Wallman, 2015).
@@
Quality of Latent Behavior Compression.  We also compared the policies encoded in the latent space of the AE with the ones used to train it. They were compared by examining the performance recovery, that is, the ratio between the performances of policies decoded from the latent space and those in the dataset. The values for MC are reported in Table 1.444Values related to the left task have been omitted as they do not present major differences from the standard task. Instead, they are reported in Table 2 of Appendix B. First, it is clear that increasing the number of latent dimensions or policy size frequently leads to better performance recovery, resulting in higher performance as well. Interestingly, some configurations appear to recover higher performances than the ones in the training dataset. This may be due to the generalization abilities of the AE, but it may also be influenced by variance in the policy evaluation process. On the contrary, we note that 1D latent spaces trained on Small policies fail to learn any meaningful encoding of the behaviors, collapsing to a uniform representation. We attribute this phenomenon to the instability of the learning process when the latent dimensions are not sufficient. Interestingly, this issue is almost always fixed by increasing the number of latent dimensions and does not arise with large policies, which show extremely good performance recovery. Finally, our analysis does not indicate that the dimension of the dataset has any meaningful influence on performance recovery.
@@
Table 1: Quality of Latent Behavior Compression in MC. We report the performance recovery for three tasks. We report mean and standard deviation computed over 3 seeds.
@@
Small
10k $0.51_{{\pm.00}}$ $0.66_{{\pm.11}}$ $0.74_{{\pm.16}}$ $0.15_{{\pm.05}}$ $0.15_{{\pm.08}}$ $0.27_{{\pm.06}}$ $0.16_{{\pm.10}}$ $0.16_{{\pm.09}}$ $0.27_{{\pm.05}}$
50k $0.64_{{\pm.19}}$ $0.93_{{\pm.10}}$ $0.94_{{\pm.06}}$ $0.10_{{\pm.10}}$ $0.42_{{\pm.15}}$ $0.44_{{\pm.20}}$ $0.11_{{\pm.11}}$ $0.45_{{\pm.14}}$ $0.47_{{\pm.28}}$
100k $0.50_{{\pm.00}}$ $0.72_{{\pm.21}}$ $0.72_{{\pm.21}}$ $0.15_{{\pm.01}}$ $0.40_{{\pm.37}}$ $0.32_{{\pm.14}}$ $0.29_{{\pm.04}}$ $0.40_{{\pm.23}}$ $0.42_{{\pm.16}}$
Medium
10k $0.83_{{\pm.23}}$ $1.01_{{\pm.01}}$ $1.02_{{\pm.00}}$ $0.25_{{\pm.12}}$ $0.84_{{\pm.02}}$ $0.84_{{\pm.10}}$ $0.36_{{\pm.22}}$ $0.71_{{\pm.20}}$ $0.78_{{\pm.24}}$
50k $0.66_{{\pm.21}}$ $1.01_{{\pm.01}}$ $1.02_{{\pm.00}}$ $0.14_{{\pm.04}}$ $0.85_{{\pm.05}}$ $0.93_{{\pm.07}}$ $0.15_{{\pm.04}}$ $0.45_{{\pm.03}}$ $0.47_{{\pm.04}}$
100k $0.51_{{\pm.00}}$ $1.02_{{\pm.00}}$ $1.02_{{\pm.00}}$ $0.14_{{\pm.03}}$ $0.60_{{\pm.24}}$ $0.97_{{\pm.02}}$ $0.22_{{\pm.04}}$ $0.44_{{\pm.01}}$ $0.53_{{\pm.10}}$
Large
10k $2_{{\pm.00}} 1.01±.0$$1_{{\pm.00}}$ $1.01_{{\pm.00}}$ $0.79_{{\pm.22}}$ $0.87_{{\pm.06}}$ $1.04_{{\pm.02}}$ $0.78_{{\pm.14}}$ $0.84_{{\pm.07}}$ $0.87_{{\pm.06}}$
50k $2_{{\pm.00}} 1.01±.0$$1_{{\pm.00}}$ $1.01_{{\pm.00}}$ $0.68_{{\pm.06}}$ $0.97_{{\pm.05}}$ $1.00_{{\pm.01}}$ $0.47_{{\pm.03}}$ $0.55_{{\pm.11}}$ $0.57_{{\pm.13}}$
100k $1_{{\pm.00}} 1.01±.0$$1_{{\pm.00}}$ $1.01_{{\pm.00}}$ $0.73_{{\pm.15}}$ $2_{{\pm.06}} 0.99±.0$$9_{{\pm.01}}$ $0.39_{{\pm.07}}$ $0.54_{{\pm.04}}$ $0.74_{{\pm.27}}$
Takeaways.  With these experiments, we provided a positive answer to (Q1): the proposed unsupervised pipeline is indeed capable of encoding behaviorally meaningful policies in a wide range of configuration and in multiple environments, ultimately leading to a compression of up to five orders of magnitude  [^5] More precisely, a 121801:1 compression rate at peak. As for (Q2), we found that while larger policies produce richer behavioral manifolds, even a one-dimensional latent space is often sufficient to capture a wide range of behaviors, supporting the hypothesis that the intrinsic dimensionality of the policy behavior manifold is dictated by the environment complexity rather than by the cardinality of the parametrization. Finally, we extracted some evidence for the existence of a critical intrinsic dimension in the behavioral manifold, but how to leverage this evidence to learn the best latent representation possible is out of the scope of the present work.

### v2

@@ 4 Method: Unsupervised Compression of the Policy Space
To address the sample inefficiency inherent in high-dimensional policy parameter spaces, we propose a paradigm shift from directly optimizing in the parameter space to learning within a compact, low-dimensional policy manifold that captures the true diversity of behaviors. This is achieved through a two-phase framework: a completely unsupervised, task-agnostic pre-training phase to discover the manifold, followed by a supervised, task-specific fine-tuning phase. As illustrated in Fig. 2, this framework is composed of three steps: $\mathbf{(1)}$ generating a behaviorally diverse dataset of policies, $\mathbf{(2)}$ learning the latent policy manifold via a generative model, and $\mathbf{(3)}$ performing fine-tuning by optimizing over this learned latent space.
Policy Dataset Generation.  Many manifold reconstruction algorithms depend on efficiently covering the manifold with samples (Bernstein et al., 2000; Cheng et al., 2005; Fefferman et al., 2016). Thus, the first stage of our framework involves generating a dataset ${\mathcal{D}}_{\Theta}$ of policies intended to cover the manifold of behaviorally diverse policies.
A naïve option is to randomly sample $N$ policies, $\hat{\mathcal{D}}_{\Theta}=\{{\bm{\theta}}_{i}\}_{i=1}^{N}$, by drawing their parameters from a uniform distribution. Unfortunately, it is well-understood that such naïve sampling from the parameter space is unlikely to produce uniform coverage of the behavior space, as it tends to favor functionally similar, often non-exploring, policies.
To address this bias, an explicit measure of behavioral diversity is needed. Looking at Eq. 4, one notices that optimizing such a measure directly requires estimating the divergence between two state-action distributions $d_{\pi_{\bm{\theta}}}^{sa},d_{\pi_{{\bm{\theta}}^{\prime}}}^{sa}$. Unfortunately, this would not only be computationally intensive but also require sampling from the environment for a potentially vast set of policies. To avoid this, we will take into account an upper bound to this quantity in the case of finite-horizon tasks (Prop. E.1, Metelli et al., 2018), namely $D(\pi_{\bm{\theta}}\parallel\pi_{{\bm{\theta}}^{\prime}})$. In practice, we substitute this measure with the L2 distance of two policies in the action space, evaluated on a finite subset of the state space, or formally:
$D(\pi_{\bm{\theta}}\parallel\pi_{{\bm{\theta}}^{\prime}})=\sqrt{\sum_{i=1}^{M}\left(\pi_{\bm{\theta}}\left(\cdot\;\middle|\;s_{i}\right)-\pi_{{\bm{\theta}}^{\prime}}\left(\cdot\;\middle|\;s_{i}\right)\right)^{2}}.$
Based on this proxy, we apply a Novelty Search algorithm (Lehman & Stanley, 2011) by computing a novelty score, $\rho(\pi_{\bm{\theta}})$, for each policy based on its average divergence from its $k_{n}$-nearest neighbors: $\rho(\pi_{\bm{\theta}})=\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}D(\pi_{\bm{\theta}}\parallel\pi_{{\bm{\theta}}_{i}}).$
@@
Latent Behavior Compression.  In the second stage, we learn the low-dimensional manifold from the filtered policy dataset ${\mathcal{D}}_{\Theta}$. Potentially, any generative model would do the work. Still, here we are interested in learning latent low-dimensional representations while preserving the end-to-end differentiability that makes gradient-based optimization effective. For these reasons, we employ a symmetric autoencoder architecture with an encoder $f_{\bm{\xi}}:\Theta\to{\mathcal{Z}}$ and a decoder $g_{\bm{\zeta}}:{\mathcal{Z}}\to\Theta$. While a standard autoencoder minimizes a parameter reconstruction error, our goal is to compress policy behavior. We therefore introduce a novel Behavioral Reconstruction Loss, which trains the autoencoder to minimize the expected behavioral divergence between the original policy and its reconstruction:
${\mathcal{L}}_{B}({\bm{\xi}},{\bm{\zeta}})=\mathbb{E}_{{\bm{\theta}}\sim{\mathcal{D}}_{\Theta}}\left[D\left(\pi_{\bm{\theta}}\parallel\pi_{(g_{\bm{\zeta}}\circ f_{\bm{\xi}})({\bm{\theta}})}\right)\right].$
This objective frees the decoder from reproducing the exact parameter values, allowing it to discover any parameterization that generates the desired behavior. As a result, the latent space $\mathcal{Z}$} becomes organized purely by functional similarity, effectively capturing the policy manifold. In practice, we use an empirical estimator of the behavioral reconstruction loss based on the notion of divergence in the action space. For this purpose, we train our autoencoders to minimize the Mean Squared Error between action vectors relative to a subset of the state space sampled at each gradient step, resulting in the estimator $\hat{\mathcal{L}}_{B}({\bm{\xi}},{\bm{\zeta}})=\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\left\|\pi_{{\bm{\theta}}_{i}}(s_{j})-\pi_{(g_{\bm{\zeta}}\circ f_{\bm{\xi}})({\bm{\theta}}_{i})}(s_{j})\right\|_{2}^{2}$, where $N$ is the number of policies, and $M$ is the number of sampled states.
@@
$\nabla_{\bm{z}}J^{R}({\bm{z}})=\nabla_{\bm{z}}g_{{\bm{\zeta}}^{\star}}({\bm{z}})^{\top}\nabla_{\bm{\theta}}J^{R}(\theta),$ (5)
@@
Remarks.  In this section, we proposed three specific instantiations for each phase. Yet, we emphasize that the proposed pipeline represents the most relevant contribution per se, independently of how it is realized, i.e., how the policies are collected, which divergence measure is used, which generative model or PO algorithm over the latent space is employed.
@@ 5 Experiments
We now investigate through extensive empirical corroboration how the proposed method addresses the research questions. In order to do so, we will mainly focus on the unsupervised pre-training phase of the proposed pipeline seen in Fig. 2, in which a latent representation is built out of general datasets of policies not designed to address any specific task explicitly, and we report the empirical results in Subsec. 5.1. Finally, we make sure that such a latent space can indeed be leveraged in later supervised fine-tuning phases as soon as a task is provided, and report the results in Subsec. 5.2. A detailed description of the environments and experimental settings can be found in Appendix B.
Experimental Domains.  The experiments are performed to illustrate essential features of Latent Behavior Compression, and for this reason, the domains are selected for being challenging while keeping high interpretability. The first is the Mountain Car Continuous (MC, Moore, 1990) environment. To evaluate the quality and characteristics of the latent space, we define four downstream tasks: standard and left have the goal state on the right and left hill, respectively; speed and height incentivize the car to keep a high speed and vertical coordinate, respectively, without terminating the episode. We also consider three environments from the MuJoCo suite (Todorov et al., 2012; Tassa et al., 2018). For Reacher (RC), we define four downstream tasks: speed, which incentivizes the fingertip to move with high linear velocity; clockwise and c-clockwise reward the agent for each step the fingertip is rotating clockwise and counterclockwise, respectively; and radial, which promotes the retraction and extension of the arm. For Hopper (HP), we define four downstream tasks: forward, backward, and standstill reward the agent for positive, negative, or close-to-zero velocity along the $x$-axis respectively; jump rewards the agent for achieving a certain position along the $z$-axis. Finally, for HalfCheetah (HC), we define four downstream tasks: forward and backward are defined as for HP; frontflip and backflip reward the agent each time it performs a frontflip and backflip, respectively.
Experimental Regimes.  The experiments are performed over a set of different parameters. In MC, we took into account three Policy Sizes (Small, Medium, and Large) with roughly $10^{1},10^{3}$, and $10^{5}$ parameters respectively, three Policy Dataset Sizes (10k, 50k, and 100k generated policies, with a 10% novelty-based cut-down), and three Latent Space Sizes (1D, 2D, and 3D). In RC, we focused on a specific configuration with Medium policies, Policy Dataset size of 100k, but five possible latent space sizes (1D, 2D, 3D, 5D, and 8D). In HP and HC, we focused on Policy Datasets of 10k, using Medium policies, with three latent space sizes (5D, 8D, and 16D). [^3] Notably, the AE architecture and training hyper-parameters are left the same for every experiment, regardless of configuration or environment.
@@ 5.1 Unsupervised Pre-training
(Q1) Is it possible to learn a low-dimensional latent representation of a high-dimensional policy parameter space in an unsupervised fashion? (Q2) What are the properties of such a latent representation, if it exists? Is its intrinsic dimension a function of the behavioral complexity rather than the size of the parameter space?
@@
Landscape of the Latent Behavior Manifold.  We visually inspected the latent spaces trained under different conditions and environments, and we report a handful in Fig. 3. Interestingly, it is apparent that the latent spaces, regardless of the choice of encoding dimension (top-to-bottom) or policy size (left-to-right), can encode some behaviorally diverse policies with high performance. For instance, in Fig. 3(e),3(f), and 3(g), a 2D latent space can encode policies of all three sizes, but the landscape grows more complex with the larger policy sizes. We speculate that this is due to the increased range of behaviors expressed by larger policies and the hardness of high-compression regimes. Indeed, the same trend is present for different tasks, as in Fig. 3(i),3(k). On the other hand, by changing the encoding dimension as in Fig. 3(b),3(f),3(j), it is clear how certain behavioral areas at high performance are able to grow larger, creating a better optimization landscape. Unfortunately yet, the compression is only as good as the dataset used to learn the latent space: when a behavior is scarcely represented in the dataset, as is the case for the task height in Fig. 3(a),3(c), it is unlikely that the learned representation will encode it in large areas, or encode it at all. As for RC, the compression architecture struggles to compress the policies at higher compression regimes (Fig. 3(d),3(h)), as the environment is more challenging and presents a wider range of behaviors. On the other hand, large areas of good quality compression are present for larger dimensions of the encoding (Fig. 3(l)), confirming the expected theoretical behavior (Hurewicz & Wallman, 2015).
@@
Quality of Latent Behavior Compression.  We also compared the policies encoded in the latent space with the ones in the training dataset. They were compared by examining the performance recovery, that is, the ratio between the performances of policies decoded from the latent space and those in the dataset. The values for MC are reported in Table 1.444Values related to the left task have been omitted as they do not present major differences from the standard task. Instead, they are reported in Table 3 of Appendix B. First, it is clear that increasing the number of latent dimensions or policy size frequently leads to better performance recovery, resulting in higher performance as well. Interestingly, some configurations appear to recover higher performances than the ones in the training dataset. This may be due to the generalization abilities of the AE, but it may also be influenced by variance in the policy evaluation process. On the contrary, we note that 1D latent spaces trained on Small policies fail to learn any meaningful encoding of the behaviors, collapsing to a uniform representation. We attribute this phenomenon to the instability of the learning process when the latent dimensions are not sufficient. Interestingly, this issue is almost always fixed by increasing the number of latent dimensions and does not arise with large policies, which show excellent performance recovery. Finally, our analysis does not indicate that the dataset’s dimension has any meaningful influence on performance recovery.
@@
Table 1: Quality of Latent Behavior Compression in MC. We report the performance recovery for three tasks. We report mean and standard deviation computed over 3 seeds. Bold indicates best across latent dimensions for each experiment setting; shading indicates best per task.
@@
Small 10k $0.51_{{\pm.00}}$ $0.66_{{\pm.11}}$ $\mathbf{0.74}_{{\pm.16}}$ $0.15_{{\pm.05}}$ $0.15_{{\pm.08}}$ $\mathbf{0.27}_{{\pm.06}}$ $0.16_{{\pm.10}}$ $0.16_{{\pm.09}}$ $\mathbf{0.27}_{{\pm.05}}$
50k $0.64_{{\pm.19}}$ $0.93_{{\pm.10}}$ $\mathbf{0.94}_{{\pm.06}}$ $0.10_{{\pm.10}}$ $0.42_{{\pm.15}}$ $\mathbf{0.44}_{{\pm.20}}$ $0.11_{{\pm.11}}$ $0.45_{{\pm.14}}$ $\mathbf{0.47}_{{\pm.28}}$
100k $0.50_{{\pm.00}}$ $\mathbf{0.72}_{{\pm.21}}$ $\mathbf{0.72}_{{\pm.21}}$ $0.15_{{\pm.01}}$ $\mathbf{0.40}_{{\pm.37}}$ $0.32_{{\pm.14}}$ $0.29_{{\pm.04}}$ $0.40_{{\pm.23}}$ $\mathbf{0.42}_{{\pm.16}}$
[…]
Generalization.  We further investigated the generalization capabilities of Latent Behavior Compression in more complex environments. Focusing on larger latent spaces where grid visualization is infeasible, we estimate performance recovery via random sampling of the latent space instead. Table 2 reports the results for HC and HP. These experiments corroborate the observations in MC and provide statistically significant evidence that the latent space generalizes beyond the training set. More specifically, increasing the latent dimension consistently improves performance recovery in HC, while such trend is likely obscured in HP by the high variance in the sampling process. Finally, we observe that generalization varies by task complexity, as indicated by the lower recovery rates for difficult tasks like frontflip and backflip in HC.
Table 2: Quality of Latent Behavior Compression in HP and HC. We report the performance recovery for four tasks by the mean performance recovery and 95% confidence interval over 10 seeds. Bold indicates best per task.
Environment Task Latent Dimensions
5D 8D 16D
Mean 95% CI Mean 95% CI Mean 95% CI
Hopper forward 1.33 $[1.03,1.69]$ 1.59 $[1.21,1.99]$ 1.48 $[1.21,1.78]$
backward 2.66 $[0.95,5.96]$ 1.29 $[1.11,1.47]$ 1.20 $[1.07,1.33]$
standstill 1.49 $[1.06,2.08]$ 1.51 $[1.02,2.25]$ 1.57 $[1.07,2.32]$
jump 3.83 $[0.64,9.32]$ 1.54 $[1.07,2.14]$ 2.42 $[1.35,3.75]$
HalfCheetah forward 1.63 $[1.02,2.31]$ 1.80 $[1.20,2.47]$ 1.84 $[1.41,2.39]$
backward 1.27 $[1.04,1.55]$ 1.52 $[1.19,1.85]$ 1.72 $[1.37,2.06]$
frontflip 0.54 $[0.29,0.86]$ 0.74 $[0.49,0.99]$ 1.20 $[0.80,1.63]$
backflip 0.55 $[0.31,0.86]$ 0.75 $[0.51,0.99]$ 1.23 $[0.81,1.66]$
Takeaways.  With these experiments, we provided a positive answer to (Q1): the proposed unsupervised pipeline is indeed capable of encoding behaviorally meaningful policies in a wide range of configuration and in multiple environments, ultimately leading to a compression of up to five orders of magnitude. [^5] More precisely, a 121801:1 compression rate at peak. As for (Q2), we found that while larger policies produce richer behavioral manifolds, even a one-dimensional latent space is often sufficient to capture a wide range of behaviors, supporting the hypothesis that the intrinsic dimensionality of the policy behavior manifold is dictated by the environment complexity rather than by the cardinality of the parameterization. Additionally, we study the scalability and generalization capabilities of the latent space. Finally, we extracted some evidence for the existence of a critical intrinsic dimension in the behavioral manifold, but how to leverage this evidence to learn the best latent representation possible is out of the scope of our work.